OpenStudy (anonymous):

Find: Lim x->0 (x^2 * cos(1/x^2))

OpenStudy (anonymous):

I know that x^2 will make f(x) = 0 but when you find the limit of cos(1/x^2), thats cos(infinity) which should be undefined. In that case would the answer be there is no limit?

OpenStudy (zarkon):

squeeze theorem

OpenStudy (zarkon):

\[-1\le \cos(x)\le1\] \[-1\le\cos\left(\frac{1}{x^2}\right)\le 1\] \[-x^2\le x^2\cos\left(\frac{1}{x^2}\right)\le x^2\]

OpenStudy (dumbcow):

Limit is 0